The perimeter of a rectangle is 120ft. The length is six more than twice the width. Find the length
Answer by Cromlix(4381)

Hi there,
Make the width = x
Length = 6 + 2x
Perimeter = 2 x Width + 2 x Length
120 = 2(x) + 2(6 + 2x)
120 = 2x + 12 + 4x
collect like terms:
2x + 4x = 120 - 12
6x = 108
x = 18
Width = 18 feet
Length = 42 feet.
Hope this helps :-)